OpenStudy (anonymous):

A triangle has sides of lengths 12, 14, and 19. Is it a right triangle? Explain.

OpenStudy (johnweldon1993):

Check using the Pythagorean Theorem...A² + B² = C² C is the hypotenuses...which would be the biggest number here so check so 12² + 14² = 19²? written differently does \[\sqrt{12² + 14²} = 19?\] No, not quite....19² is 361 12² = 144 and 14² is 196...196+144 is 340 340 doesnt equal 361 lol so no it is NOT a right triangle